# Unveiling dynamical quantum error correcting codes via non-invertible symmetries

### v1

@@ I Introduction
Quantum error correction is essential for building a fault-tolerant quantum computer. Among the most studied constructions are quantum stabilizer codes Calderbank97 ; Gottesman:1997zz ; Gottesman:1998se . Recent work has emphasized the role of dynamical stabilizer codes (DSCs), obtained by evolving a stabilizer code through a sequence of measurements hastings2021dynamically ; Haah2022boundarieshoneycomb ; Aasen22 ; Gidney2021faulttolerant ; Davydova23 ; Kesselring:2022eax ; Zhang23 ; Bombin2024 ; Gidney2023 ; Bauer2024topologicalerror ; Ellison23 ; Davydova:2023mnz ; Dua24 ; Townsend-Teague23 ; Higgott24 ; GransSamuelsson2024improvedpairwise ; Fahimniya25 ; Bauer2025lowoverheadnon ; Bauer2025x+y ; MdlF25 ; Xu:2025zos . In a DSC, the code subspace depends on the measurement sequence, and these codes exhibit several desirable features: stabilizers with low-weight Pauli operators and the ability to natively and fault-tolerantly implement logical operations. However, their dynamical nature makes error detection and correction more challenging. Unlike static stabilizer codes, one must adopt a spacetime perspective of errors, keeping track of both their spatial location and temporal occurrence Delfosse:2023aqk ; Fu:2024hin ; Blackwell25 .
@@
In this paper, we given an answer to this question by establishing a one-to-one correspondence between qudit stabilizer codes and non-anomalous 2-form symmetries of a class of 4+1d TQFTs  [^1] The primary operators of certain 1+1d RCFTs and the corresponding line operators of their associated 3d Chern-Simons theories capture the properties of some static qudit stabilizer codes. However, this setting is limited to special stabilizer codes. This boils down to the fact that the braiding of line operators in 2+1d is symmetric, while the abelianized Pauli group has a symplectic structure Dymarsky:2020qom ; Dymarsky:2021xfc ; Buican:2021uyp ; Barbar:2023ncl ; Buican:2023ehi .. These symmetries are implemented by topological 2-dimensional surface operators in the TQFT. Under this correspondence, the measurement of a Pauli operator is realized as a 4-dimensional topological operator. The effect of a measurement on the stabilizer and logical operators of the code can be obtained from the action of the corresponding 4-dimensional operator on topological surface operators. The irreversibility of a measurement is directly related to the fact that these 4-dimensional operators implement non-invertible symmetries of the TQFT.
@@ II.3 Detectors of DSC
$\mathcal{I}_{B}(U)=\begin{cases}\sum_{U^{\prime}\in B}U^{\prime}\times U~~&\text{ if }\beta(U,U^{\prime})=1,\\ 0&\text{ otherwise}.\end{cases}$ (18)
@@ IV.1 Measurements and non-invertible symmetries
A subset of Pauli operators $\mathcal{M}$ which commute with each other can be simultaneously measured. In fact, if commuting operators $M_{1}$ and $M_{2}$ are measured, the measurement outcome of $M_{1}M_{2}$ is also determined. Therefore, we can take a commuting subgroup $\mathcal{M}$ of $\mathcal{P}_{n}$ whose generators can be simultaneously measured. Under the isomorphism $\phi$, these are in one-to-one correspondence with non-anomalous subgroups of $A$. We can use this correspondence to study DSCs in terms of the 2-form gauge theory. To that end, let us first consider a static stabilizer code defined on $n$ qubits, which can be regarded as a DSC with measurement sequence $\mathcal{M}_{t}=\mathcal{S},\forall t>0$ where $\mathcal{S}$ is a abelian subgroup of the Pauli group. This defines a static stabilizer code with stabilizer group $\mathcal{S}$. Under the isomorphism $\phi$, $\phi(\mathcal{S})$ is a non-anomalous subgroup of $A$. Conversely, each non-anomalous subgroup of $A$ corresponds to a static stabilizer code. Moreover, any 2-form gauge theory admits a gapped boundary obtained by gauging a maximal non-anomalous subgroup of A. When the 2-form gauge theory is interpreted as the SymTFT of a 4-dimensional QFT with a fixed symmetry Apruzzi:2021nmk ; Chatterjee:2022kxb ; Freed:2022qnc ; Kaidi:2022cpf , its different gapped boundaries correspond to the different possible ways of gauging the symmetry of the 4d theory. Through the isomorphism $\phi$, these gapped boundaries are in one-to-one correspondence with stabilizer codes that have a 1-dimensional code subspace.
@@ VI Acknowledgments
RR would like to thank Matthew Buican, Clement Delcamp, and Anatoly Dymarsky for discussions related to this project. This work was initiated during a visit to the Simons Center for Geometry and Physics, whose hospitality is gratefully acknowledged. RR also thanks the International Centre for Mathematical Sciences for their hospitality during a Research-in-Groups programme.

### v2

@@ I Introduction
Quantum error correction is essential for building a fault-tolerant quantum computer. Among the most studied constructions are quantum stabilizer codes Calderbank97 ; Gottesman:1997zz ; Gottesman:1998se . Recent work has emphasized the role of dynamical stabilizer codes (DSCs), obtained by evolving a stabilizer code through a sequence of measurements hastings2021dynamically ; Haah2022boundarieshoneycomb ; Aasen22 ; Gidney2021faulttolerant ; Davydova23 ; Kesselring:2022eax ; Zhang23 ; Bombin2024 ; Gidney2023 ; Bauer2024topologicalerror ; Ellison23 ; Davydova:2023mnz ; Dua24 ; Townsend-Teague23 ; Kobayashi:2023zxs ; Higgott24 ; GransSamuelsson2024improvedpairwise ; Fahimniya25 ; Bauer2025lowoverheadnon ; Bauer2025x+y ; MdlF25 ; Xu:2025zos . In a DSC, the code subspace depends on the measurement sequence, and these codes exhibit several desirable features: stabilizers with low-weight Pauli operators and the ability to natively and fault-tolerantly implement logical operations. However, their dynamical nature makes error detection and correction more challenging. Unlike static stabilizer codes, one must adopt a spacetime perspective of errors, keeping track of both their spatial location and temporal occurrence Delfosse:2023aqk ; Fu:2024hin ; Blackwell25 .
@@
In this paper, we given an answer to this question by establishing a one-to-one correspondence between qudit stabilizer codes and non-anomalous 2-form symmetries of a class of 4+1d TQFTs  [^1] The primary operators of certain 1+1d RCFTs and the corresponding line operators of their associated 3d Chern-Simons theories capture the properties of some static qudit stabilizer codes. However, this setting is limited to special stabilizer codes. This boils down to the fact that the braiding of line operators in 2+1d is symmetric, while the abelianized Pauli group has a symplectic structure Dymarsky:2020qom ; Dymarsky:2021xfc ; Buican:2021uyp ; Buican:2023ehi .. These symmetries are implemented by topological 2-dimensional surface operators in the TQFT. Under this correspondence, the measurement of a Pauli operator is realized as a 4-dimensional topological operator. The effect of a measurement on the stabilizer and logical operators of the code can be obtained from the action of the corresponding 4-dimensional operator on topological surface operators. The irreversibility of a measurement is directly related to the fact that these 4-dimensional operators implement non-invertible symmetries of the TQFT.
@@ II.3 Detectors of DSC
$\mathcal{I}_{B}(U)=\begin{cases}\sum_{U^{\prime}\in B}U^{\prime}\times U~~&\text{ if }\beta(U,U^{\prime})=1~\forall~U^{\prime}\in B,\\ 0&\text{ otherwise}.\end{cases}$ (18)
@@ IV.1 Measurements and non-invertible symmetries
A subset of Pauli operators $\mathcal{M}$ which commute with each other can be simultaneously measured. In fact, if commuting operators $M_{1}$ and $M_{2}$ are measured, the measurement outcome of $M_{1}M_{2}$ is also determined. Therefore, we can take a commuting subgroup $\mathcal{M}$ of $\mathcal{P}_{n}$ whose generators can be simultaneously measured. Under the isomorphism $\phi$, these are in one-to-one correspondence with non-anomalous subgroups of $A$. We can use this correspondence to study DSCs in terms of the 2-form gauge theory. To that end, let us first consider a static stabilizer code defined on $n$ qubits, which can be regarded as a DSC with measurement sequence $\mathcal{M}_{t}=\mathcal{S},\forall t>0$ where $\mathcal{S}$ is a abelian subgroup of the Pauli group. This defines a static stabilizer code with stabilizer group $\mathcal{S}$. Under the isomorphism $\phi$, $\phi(\mathcal{S})$ is a non-anomalous subgroup of $A$. Conversely, each non-anomalous subgroup of $A$ corresponds to a static stabilizer code. Moreover, any 2-form gauge theory admits a gapped boundary obtained by gauging a maximal non-anomalous subgroup of A. When the 2-form gauge theory is interpreted as the SymTFT of a 4-dimensional QFT with a fixed symmetry Gaiotto:2020iye ; Apruzzi:2021nmk ; Chatterjee:2022kxb ; Freed:2022qnc ; Kaidi:2022cpf , its different gapped boundaries correspond to the different possible ways of gauging the symmetry of the 4d theory. Through the isomorphism $\phi$, these gapped boundaries are in one-to-one correspondence with stabilizer codes that have a 1-dimensional code subspace.
@@ VI Acknowledgments
RR would like to thank Matthew Buican, Clement Delcamp, and Anatoly Dymarsky for discussions related to this project. This work was initiated during a visit to the Simons Center for Geometry and Physics, whose hospitality is gratefully acknowledged. RR also thanks the International Centre for Mathematical Sciences for their hospitality during a Research-in-Groups programme. The work of RR is supported by the UKRI Frontier Research Grant, underwriting the ERC Advanced Grant “Generalized Symmetries in Quantum Field Theory and Quantum Gravity”.